### v4

# Galactic Trajectories of Interstellar Objects 1I/’Oumuamua, 2I/Borisov, and 3I/Atlas

Abstract
Context. The first interstellar objects, 1I/‘Oumuamua, 2I/Borisov and 3I/ATLAS, were discovered over the past decade.
Aims. We follow the trajectories of known interstellar objects in the gravitational potential of the Milky Way galaxy to constrain their possible origin.
Methods. We initiate the trajectories based on the measured velocities of the interstellar objects relative to the Local Standard of Rest. Since the scale-height of stars in the Milky-Way disk increases with age, we use the vertical excursion of each interstellar object from the Milky-Way disk mid-plane to constrain their likely age.
Results. The small vertical extent of 1I/‘Oumuamua’s past trajectory suggests that it originated near the mid-plane of the thin Galactic disk of stars, implying a likely age younger than 1–2 Gyr. The maximal excursion of the comet 2I/Borisov is similar to that of the Sun, suggesting a similar age. The interstellar object 3I/ATLAS exhibits significantly larger vertical excursions, with $z_{\text{max}}=0.475$ , suggesting that it originated from an older source population in the thick Galactic disk.
Key Words.:
interstellar objects – galactic trajectories – origins
1 Introduction
The discovery of interstellar objects over the past decade has sparked significant interest in understanding their origins and dynamics (see reviews by Siraj & Loeb (2022a), Jewitt and Seligman (2023), Jewitt (2024), and references therein). A fundamental unknown is the likely origin of each of these objects (Bailer-Jones et al., 2018, 2020). Constraining the sources could shed light on the nature of these interstellar objects and the astrophysical processes that created them (see, for example, Zhang & Lin (2020) or Loeb & MacLeod (2024)).
In this paper, we numerically integrate the trajectories of interstellar objects back in time in the gravitational potential of the Milky-Way as a way to relate them to potential stellar populations. For simplicity, we ignore transient gravitational features such as spiral arms and the Galactic bar. This is a reasonable approximation for orbits in the outer part of the Galactic disk.
By integrating the orbits of these objects back in time, we are able to constrain the spatial region of their potential sources within the Milky Way. These constraints limit the possiblr birthplaces of the different interstellar objects and provide insights into the Galactic environment from which they originated.
Since the scale-height of stars in the Milky-Way disk increases with age, we use the vertical excursion of each interstellar object from the Milky-Way disk mid-plane to constrain its likely age. Any dynamical pumping of the stellar scale-height by gravitational perturbations from satellite galaxies or star clusters, would affect interstellar objects and stars alike since both populations are collisionless. Hence, our constraints apply to the full age of the interstellar objects irrespective of their travel time.
The organization of this paper is as follows. Section 2 describes our method of calculation and the resulting Galactic trajectories for the individual interstellar objects, with subsections dedicated to 2I/Borisov (§2.2), 1I/‘Oumuamua (§2.3) and 3I/ATLAS (§2.4). In Section 3, we use the resulting vertical excursion of these trajectories to derive the probability distribution for the likely age for each of these objects. Finally, we summarize the implications of our results in Section 4.
2 Galactic Trajectories of Interstellar Objects
2.1 Method of Calculation
Our numerical integration is based on the OrbitIntegrator from GalPot, which utilizes the MWPotential2014 model from McMillan (2017) to simulate the gravitational potential of the Milky Way galaxy. We assume a circular velocity of 233 km s-1 and an orbital radius of 8.3 kpc for the Local Standard of Rest (LSR) around the Galactic center, consistently with the latest Gaia data (Põder et al., 2023). We initiate the trajectories from the velocity of each interstellar object relative to the LSR and follow them back in time for 10 Gyr, roughly the age of the Milky Way disk (Fantin et al., 2019). The code uses the NumPy’s linspace function that defines the relevant timescale, and then integrates the orbits using the OrbitIntegrator’s getOrbitPath and Stats method. The resulting orbit paths are extracted and plotted using Matplotlib, with separate plots showing the radial distance $R$, vertical excursion from the disk mid-plane $z$, and azimuthal angle $\phi$, as functions of time. The code also calculates the evolution of the velocity components $v_{R}$, $v_{z}$, and $v_{\phi}$ as functions of time.
In order to convert the velocity measurements in the Solar system to the Galactic frame of reference, we incorporate the motion of the Sun relative to the Local Standard of Rest (LSR), with the Galactic components: $U_{\odot}=10.79\pm 0.56$ km s-1, $V_{\odot}=11.06\pm 0.94$ km s-1, and $W_{\odot}=7.66\pm 0.43$ km s-1 (Robin et al., 2022). Before entering the solar system, 2I/Borisov’s velocity was $(U,V,W)=(33.1,-6.8,8.3)$ km s-1, while 1I/‘Oumuamua’s velocity relative to the Sun is $(U-U_{\odot},V-V_{\odot},W-W_{\odot})=(-11.457\pm 0.009,-22.395\pm 0.009,-7.74\pm 0.011)$ km s-1 (Bailer-Jones et al., 2020; Mamajek, 2017). For the interstellar object 3I/ATLAS (C/2025 N1), we adopt its inbound heliocentric velocities of $(U,V,W)=(-54.4,-20.3,+19.5)$ km s-1 and corresponding Galactocentric cylindrical velocities of $(v_{R},v_{z},v_{\phi})=(+43.62,+27.18,+223.76)$ km s-1, derived from JPL orbital solution JPL#5. We verified the orbital calculation accuracy through direct API comparison with JPL Horizons (achieving $<10^{-10}$ AU precision), independent mathematical validation of the elements-to-Cartesian conversion, and consistency checks with predicted planetary encounter geometries. The Galactocentric velocities were obtained via standard coordinate transformations incorporating solar motion parameters.
2.2 1I/’Oumuamua
The past evolution of the distance of the interstellar object 1I/‘Oumuamua from the Sun is shown in figure 1. The 1I/‘Oumuamua-Sun separation follows a period of about 2.2 Gyr, with 1I/‘Oumuamua being on the other side of the Milky Way disk relative to the Sun about 1.1 Gyr ago.
Figure 1: Distance of 1I/’Oumuamua from the Sun back in time.
The radial ($R$) and vertical ($z$) extent of 1I/‘Oumuamua’s trajectory relative to the Galactic plane are depicted in figure 2 for both 5 Gyr (left panel) and 1 Gyr (right panel) integrations. As a result of its low vertical velocity relative to the LSR, the vertical excursion of 1I/‘Oumuamua (orange) is a factor of $\sim 6$ smaller than that of the Sun (blue).
Figure 2: Trajectory of 1I/’Oumuamua in the $R-z$ plane of the Milky-Way disk. Left panel shows 5 Gyr integration, right panel shows 1 Gyr integration.
The motion in the Galactic plane ($x-y$) is shown in figure 3, comparing 5 Gyr and 1 Gyr integrations.
Figure 3: Trajectory of 1I/’Oumuamua in the $x-y$ plane of the Milky-Way disk. Left panel shows 5 Gyr integration, right panel shows 1 Gyr integration.
2.3 2I/Borisov
The past evolution of the distance of the interstellar comet 2I/Borisov from the Sun is shown in figure 4. The Sun-2I/Borisov separation follows a period of about 3.2 Gyr, with 2I/Borisov being on the other side of the Milky-Way disk relative to the Sun about 1.6 Gyr ago.
Figure 4: Distance of the comet 2I/Borisov from the Sun back in time with the present time represented by 0.
The radial ($R$) and vertical ($z$) extent of 2I/Borisov’s trajectory relative to the Galactic plane are depicted in figure 5 for both 5 Gyr (left panel) and 1 Gyr (right panel) integrations. The trajectories of 2I/Borisov (orange) resemble the corresponding ranges for the Sun (blue).
Figure 5: Trajectory of the comet 2I/Borisov in the $R-z$ plane of the Milky-Way disk. Left panel shows 5 Gyr integration, right panel shows 1 Gyr integration.
The motion in the Galactic plane ($x-y$) is shown in figure 6, again comparing 5 Gyr and 1 Gyr integrations.
Figure 6: Trajectory of the comet 2I/Borisov in the $x-y$ plane of the Milky-Way disk. Left panel shows 5 Gyr integration, right panel shows 1 Gyr integration.
2.4 3I/ATLAS
The past evolution of the distance of the interstellar object 3I/ATLAS from the Sun is shown in figure 7. The 3I/ATLAS-Sun separation follows a period of about 3.63 Gyr, with 3I/ATLAS being on the other side of the Milky Way disk relative to the Sun about 1.815 Gyr ago.
Figure 7: Distance of 3I/ATLAS from the Sun back in time.
The radial ($R$) and vertical ($z$) extent of 3I/ATLAS’s trajectory relative to the Galactic plane are depicted in figure 8 for both 5 Gyr (left panel) and 1 Gyr (right panel) integrations. As a result of its vertical velocity, the vertical excursion of 3I/ATLAS (orange) is larger than that of the Sun (blue).
Figure 8: Trajectory of 3I/ATLAS in the $R-z$ plane of the Milky-Way disk. Left panel shows 5 Gyr integration, right panel shows 1 Gyr integration.
The motion in the Galactic plane ($x-y$) is shown in figure 9, comparing 5 Gyr and 1 Gyr integrations.
Figure 9: Trajectory of 3I/ATLAS in the $x-y$ plane of the Milky-Way disk. Left panel shows 5 Gyr integration, right panel shows 1 Gyr integration.
3 Probability Distribution of Age
[…]
3.1 Star Formation Rate (SFR) as a Function of Lookback Time
We start from the empirically-calibrated star formation rate (SFR) history of the Milky-Way disk, based on the age distribution of white dwarfs (Fantin et al., 2019). The SFR is divided into two parts: a flat component from lookback time of 0 to 9 Gyr, plus a Gaussian component centered on 10 Gyr. For the flat component, we adopt a constant value of $\text{SFR}_{\text{linear}}$, whereas the Gaussian part is described by,
$\text{SFR}_{\text{Gaussian}}(t)=\text{SFR}_{\text{peak}}\times\exp\left(-{\frac{1}{2}}\left(\frac{t-\text{t}_{\text{peak}}}{\sigma}\right)^{2}\right),$ (1)
where $\text{SFR}_{\text{peak}}=3\text{SFR}_{\text{linear}}$, $\text{t}_{\text{peak}}=10$ Gyr, and $\sigma=0.5$ Gyr. Figure 10 shows the resulting age distribution of stars, plotted in 1 Gyr bins
Figure 10: Probability distribution of age, $p(t)$ in units of Gyr-1, for the stars in the Milky Way disk, based on figure 9 in Fantin et al. (2019).
The stellar mass formed in each temporal bin of width $\Delta t=$1 Gyr, is given by
$\Delta M(t)=\text{SFR}_{\text{binned}}(t)\times\Delta t.$ (2)
The current number density of disk stars at the Galactic orbital radius of the Sun follow an exponential vertical distribution relative to the mid-plane, $\propto\exp\{-|z/h||\}$, with a scale-height, $h(t)$, that depends on their age $t$. We adopt the observed $h(t)$ from Cukanovaite et al. (2023) and combine it with the star formation history described above to calibrate the fraction of stars with any given age.
To quantify the age distribution of stars within the vertical excursion of each interstellar object, we determine the number of stars residing in each age bin up to the maximum $z$ value sampled by the object’s trajectory, denoted by $z_{\text{max}}$,
$\Delta N(t)\propto\Delta M(t)\left(1-\exp\left(-\frac{z_{\text{max}}}{h(t)}\right)\right),$ (3)
where $z_{\text{max}}$ is the maximum vertical height for the object.
To obtain a unit normalization of our probability distribution of ages, we find the total number of stars within the given $z$ range as,
$N_{\text{total}}=\sum_{i}\Delta N(t_{i}),$ (4)
where $N_{\text{total}}$ represents the total number of stars and $\Delta N(t)$ denotes the number of stars in the age bin centered at age $t$. The summation is performed over all age bins $t_{i}$ given the constraint of $z_{\rm max}$.
The normalized age distribution function $p(t)$ is then calculated from the relation,
$p(t_{i})=\frac{\Delta N(t_{i})}{N_{\text{total}}\Delta t}.$ (5)
3.2 Likelihood of Ages for individual Interstellar Objects
Figure 11 shows the probability distribution of likely ages for the comet 2I/Borisov, which straddles the age of the Sun at $4.6$ Gyr.
Figure 11: Probability distribution of age, $p(t)$ in units of Gyr-1, for the interstellar comet 2I/Borisov with $z_{\text{max}}=0.12$ kpc.
In contrast, 1I/‘Oumuamua is much younger with a likely age of 1-2 Gyr, based on its much smaller value of $z_{\rm max}$, as shown in Figure 12.
Figure 12: Probability distribution of age, $p(t)$ in units of Gyr-1, for 1I/‘Oumuamua with $z_{\text{max}}=0.015$ kpc.
Finally, 3I/ATLAS may follow a probability distribution similar to the entire Galactic disk population because of its potentially large value of $z_{\rm max}$, as shown in Figure 13.
Figure 13: Probability distribution of age, $p(t)$ in units of Gyr-1, for 3I/ATLAS with $z_{\text{max}}=\text{0.475}$ kpc.
4 Conclusions
We have integrated the Galactic trajectories of known interstellar objects back in time. The vertical extent of the trajectory of 1I/’Oumuamua suggests that it originated near the disk mid-plane of the thin disk, where the ages of most stars are below 1–2 Gyr. In contrast, the comet 2I/Borisov’s vertical excursion is similar to that of the Sun, suggesting a similar age. Given its high vertical excursions, the putative source star of 3I/ATLAS is likely older than that of 1I/’Oumuamua or 2I/Borisov.
Over many orbital times, the trajectories are expected to migrate radially as a result of transient features like spiral arms or the Galactic bar (Sellwood & Binney, 2002). Our constraints apply to the full age of the interstellar objects, because they respond just like the underlying stellar population to gravitational perturbations that pump up their scale-height over time.
Our constraints on the ages of the various objects represent upper limits because the velocity dispersion of the interstellar objects includes both the velocity dispersion of their parent stars and the dispersion in their characteristic ejection speed away from their birth system.
Concluding note: As we were about to submit our manuscript for publication, Hopkins et al. (2025) posted a paper which addresses some of the same topics. Our main results are consistent with theirs.
Acknowledgements. We thank Matthew Hopkins for his careful cross-checking of our paper. This work was supported in part by Harvard University, the Black Hole Initiative and the Galileo Project.